Please please i seriously need help

The following financial statement information is from five separate companies.

Company
A
Company
B
Company
C
Company
D
Company
E
December 31, 2018
Assets $ 40,000 $ 31,200 $ 25,600 $ 71,200 $ 109,200
Liabilities 32,800 21,840 13,824 49,127 ?
December 31, 2019
Assets 45,000 32,400 ? 81,900 124,200
Liabilities ? 22,032 14,773 39,312 98,118
During year 2019
Owner investments 6,000 1,400 9,750 ? 6,500
Net income (loss) 10,550 ? 7,400 13,515 8,742
Owner cash withdrawals 3,500 2,000 5,875 0 11,000

2a. What is the amount of equity on December 31, 2018 for Company B


2b. What is the amount of equity on December 31, 2019 for Company B

2c. .Statement of OWNERS EQUITY
2.d What is net income for year 2019 for Company B
Answer #1

2a. Calculation of amount of Equity on Dec, 31 2018 for Company B

Equity = Assets - Liabilities = 31200 - 21840 = $9360

2b. Calculation of amount of Equity on Dec, 31 2019 for Company B

Equity = Assets - Liabilities = 32400 - 22032 = $10368

2c. Statement of owner's Equity on Dec, 31 2019 for Company B

Amount of Equity as on Dec,31 2018 $       9,360.00
Add;-Investment during 2019 $       1,400.00
Add;-income during 2019(Calculated under 2d.) $       1,608.00
$     12,368.00
Less;- Withdrawals during 2019 $       2,000.00
Amount of Equity as on Dec,31 2019 $     10,368.00

2d. Calculation of amount of Net income for year 2019 for Company B

Amount of Net Income = Amount of Equity as on Dec 31, 2019 + Cash Withdrawals During 2019 - Amount of Equity as on Dec 31, 2018 - Owner's Investment during 2019 year

= 10368 + 2000 - 9360 - 1400 = $1608